OpenStudy (anonymous):

]Identify the correct slope and y intercept of the equation 6x - 3y = 12.

OpenStudy (anonymous):

slope=6y-intercept=3

OpenStudy (anonymous):

That's not one of the possible answers

OpenStudy (anonymous):

intercept (Y) x=0 then F(0) = 6x0 - 3y = 12 -3y = 12 y=12/-3 = -4

OpenStudy (anonymous):

slope is 6/3 = 2 (i think)